# Improved Stochastic Optimization of LogSumExp

### v2

@@ 3.1 Continuous entropy-regularized OT
The classical optimal transport (Monge–Kantorovich) problem consists in finding a coupling of two probablility measures $\mu,\nu\in\mathcal{P}(\mathcal{X})$ which minimizes the integral of a given measurable cost function $c\colon\mathcal{X}\times\mathcal{X}\to\mathbb{R}_{+}$ (e.g., a distance), i.e., $W(\mu,\nu)\coloneqq\inf_{\pi\in\Pi(\nu,\mu)}\int c(x,z)\,\mathrm{d}\pi(x,z)$, where $\Pi(\mu,\nu)\subset\mathcal{P}(\mathcal{X}\times\mathcal{X})$ is the set of couplings (transport plans) of $\mu$ and $\nu$ (see kantorovich1942translocation; villani2008optimal; santambrogio_optimal_2015). For simplicity of demonstration, we assume that the measures are defined on the same space $\mathcal{X}$, but the results extend trivially to the case of two different spaces. Following cuturi2013sinkhorn, we consider entropy-regularized optimal transport (eOT) problem:
@@
where $\nu\otimes\mu$ is the product measure. It is known that eOT admits the following dual and semi-dual formulations (see, e.g., genevay2016stochastic):
$W_{\varepsilon}(\mu,\nu)=\underbrace{\max_{u,v\in\mathcal{C}(\mathcal{X})}\iint_{{\mathcal{X}\times\mathcal{X}}}f_{\varepsilon}(x,y,u,v)\,\mathrm{d}\mu(x)\,\mathrm{d}\nu(y)}_{\text{dual}}=\underbrace{\max_{v\in\mathcal{C}(\mathcal{X})}\int_{{\mathcal{X}}}h_{\varepsilon}(x,v)\,\mathrm{d}\mu(x)}_{\text{semi-dual}},$
@@
$f_{\varepsilon}(x,y,u,v)$ $\coloneqq u(x)+v(y)-\varepsilon\exp\left(\frac{u(x)+v(y)-c(x,y)}{\varepsilon}\right),$ (12)
$h_{\varepsilon}(x,v)$ $\coloneqq\int_{{\mathcal{X}}}v(y)\mathrm{d}\nu(y)-\varepsilon\log\left(\int_{{\mathcal{X}}}\exp\left(\frac{v(y)-c(x,y)}{\varepsilon}\right)\mathrm{d}\nu(y)\right)-\varepsilon,$ (13)
and $\varepsilon>0$ is the regularization coefficient. In the LSOT framework (seguy2018large), the potentials $u$ and $v$ are parameterized by neural networks and optimized via SGD. While Appendix B.1 contains a more detailed literature review, we briefly position LSOT among other solvers to motivate its selection as a baseline. LSOT offers two key advantages relevant to our goals: it is less computationally intensive than modern solvers requiring adversarial training (korotin2023neural; gushchin2023entropic; asadulaev2024neural) or iterative Langevin dynamics (mokrov2024energyguided), and it supports a general cost function—contrary to other efficient solvers like (korotin2024light) tailored to the quadratic cost. Therefore, to solve eOT with a general cost function under modest computational constraints, we adopt the LSOT framework as our primary baseline. In Appendix B.2 we compare also to (genevay2016stochastic) who use an RKHS parametrization for the potentials $u$ and $v$.
@@
The main drawback of this approach is the presence of the exponent in the dual objective (and consequently in the SGD updates). Specifically, exponents are prone to floating-point exceptions (goldberg1991every), especially if the regularization parameter $\varepsilon$ is relatively small, which is often the case. For example, if $\varepsilon=0.01$ and $z\geq 7.1$, then $e^{z/\varepsilon}$ exceeds the representable range of a double-precision (float64) floating-point number — an overflow happens. When single precision (float32) is used, an overflow happens even for $z\geq 0.89$.
@@
If we consider instead the semi-dual formulation and use the approximation (6), we arrive at the problem
@@
with $\tilde{h}_{\varepsilon}(x,y,v,\alpha)\coloneqq v(y)-\alpha(x)-\frac{\varepsilon}{\rho}\log\bigl(1+\rho e^{(v(y)-c(x,y)-\alpha(x))/\varepsilon}\bigr)-\varepsilon,$ (15)
which also admits neural network parameterization and optimization via SGD. One can show, in the same way as in genevay2016stochastic, that this corresponds to the regularized OT problem (11) with Safe KL divergence $D_{\rho}$ rather than the usual KL, i.e.
@@
We consider the MNIST (deng2012mnist) and EMNIST-letters (cohen2017emnist) datasets as samples from the distributions $\mu$ (digits) and $\nu$ (letters). Manhattan distance $\ell_{1}$ is chosen as the cost function for computing eOT between $\mu$ and $\nu$. We parameterize the functions $u$, $v$ in LSOT and $v$, $\alpha$ in our proposed approach using a multilayer perceptron with two hidden layers (dimensions 256 and 128) and ReLU activations. The batch size is 256, and the learning rate is selected via grid search over $\{10^{-6},10^{-5},\ldots,10^{-1}\}$. The objective is evaluated on the empirical distributions of the dedicated test sets.
@@
3.2 Distributionally robust optimization with KL divergence
One of the approaches to training a model that is robust to data distribution shifts and noisy observations is called Distributionally Robust Optimization (DRO) (kuhnDistributionallyRobustOptimization2024). In contrast to the standard Empirical Risk Minimization (ERM) approach, which minimizes the average loss on the training sample, DRO minimizes the risk for the worst-case distribution among those close to a reference measure (e.g., empirical distribution). A prominent example is KL divergence DRO (huKullbackLeiblerDivergenceConstrained), which is formulated as the saddle-point problem
@@ 3.2 Distributionally robust optimization with KL divergence
However, when $n$ is large, computing the full gradient $\nabla\mathcal{L}(\theta)=\sum_{i=1}^{n}p_{i}^{*}(\theta)\nabla\ell_{i}(\theta)$ becomes costly. A straightforward approach (levy2020large) is to sample a batch $D$, compute the respective softmax weights $p_{i}^{D}(\theta)\coloneqq\frac{e^{\ell_{i}(\theta)/\lambda}}{\sum_{j\in D}e^{\ell_{j}(\theta)/\lambda}}$, and define a gradient estimator by
@@
$\tilde{\nabla}^{D}_{\theta}G(\theta,\alpha)\coloneqq\frac{1}{|D|}\sum_{i\in D}\sigma_{\rho}\left(\textstyle\frac{\ell_{i}(\theta)-\alpha}{\lambda}\right)\nabla\ell_{i}(\theta),\quad\tilde{\nabla}^{D}_{\alpha}G(\theta,\alpha)\coloneqq 1-\frac{1}{|D|}\sum_{i\in D}\sigma_{\rho}\left(\textstyle\frac{\ell_{i}(\theta)-\alpha}{\lambda}\right).$ (20)
@@
Consider the California housing dataset (pace1997sparse) consisting of 20,640 objects represented by 8 features. Let $\ell_{i}$ be the squared error of a linear model, $\ell_{i}(\theta)=(y_{i}-\theta^{\top}x_{i})^{2}$. We use accelerated SGD with the gradient estimator (18) (levy2020large) as the baseline approach for solving (17), and compare it to our proposed gradient estimator (20). We consider various values of the penalty coefficient $\lambda\in\{1/5,1,5\}$ and batch sizes $|D|\in\{10,10^{2},10^{3}\}$. For each configuration, we select the optimal learning rate from $\{10^{-9},10^{-8},\ldots,10^{-4}\}$. The approximation accuracy parameter $\rho$ in our method is varied across $\{10^{-1},10^{-3},10^{-5}\}$. Momentum is fixed at 0.9 (without tuning), and the least squares solution is used as the initial point for optimization.
Table 1: Objective value (17) (mean $\pm$ std across 10 runs) at epoch 50 for baseline (18) (levy2020large) and proposed gradient estimator (20) with different $\rho$ values. Results are shown for various penalty coefficients $\lambda$ and batch sizes $|D|$, with optimal learning rates selected from $\{10^{-9},\dots,10^{-4}\}$. Best results per column are shown in bold.
$\lambda=1/5$ $\lambda=1$ $\lambda=5$
@@
3.3 Distributionally robust optimization with unbalanced OT
In the KL divergence DRO described in the previous subsection, uncertainty set is limited to distributions with the same support as the empirical measure $\mu=\frac{1}{n}\sum_{i}\delta_{x_{i}}$. Another popular approach, Wasserstein DRO (WDRO) (mohajerinesfahaniDatadrivenDistributionallyRobust2018; sinhaCertifyingDistributionalRobustness2020), considers the worst-case risk over shifts within a Wasserstein (OT) ball around a reference measure $\mu$ instead of the KL-ball in (16), thus including continuous probability measures. Unfortunately, this approach is not resilient to outliers that are geometrically far from the clean distribution since OT metric is sensitive to them (nietert2023outlier). A natural generalization is to switch to semi-balanced OT (liero_optimal_2018; chizat_unbalanced_2019; kondratyevNewOptimalTransport2016), which replaces a hard constraint on one of the marginals with a mismatch penalty function, e.g.,
$W_{\beta}(\nu,\mu)=\inf_{\begin{subarray}{c}\pi\in\mathcal{P}(\mathcal{X}\times\mathcal{X})\\ \pi_{1}=\nu\end{subarray}}\int_{{\mathcal{X}\times\mathcal{X}}}c(x,z)\,\mathrm{d}\pi(x,z)+\beta\,D_{KL}(\pi_{2},\mu),$
@@ 3.3 Distributionally robust optimization with unbalanced OT
where $\lambda>0$ is the Lagrangian penalty parameter. Using standard duality, wang2024outlier showed that when $\mu=\frac{1}{n}\sum_{i}\delta_{x_{i}}$ is the empirical distribution, this is equivalent to
$\min_{\theta\in\Theta}F(\theta)\coloneqq\lambda\beta\,\log\Bigl(\frac{1}{n}\sum_{i=1}^{n}e^{\hat{\ell}_{i}(\theta)/(\lambda\beta)}\Bigr)\quad\text{with }\;\hat{\ell}_{i}(\theta)\coloneqq\sup_{z\in\mathcal{X}}\{\ell(\theta;z)-\lambda c(z,x_{i})\},$ (21)
To avoid the costly gradient computation of LogSumExp, the authors drop the logarithm and use SGD to optimize the sum of exponents,
$\min_{\theta\in\Theta}\frac{1}{n}\sum_{i=1}^{n}e^{\hat{\ell}_{i}(\theta)/(\lambda\beta)}.$ (22)
@@
$\min_{\begin{subarray}{c}\theta\in\Theta\\ \alpha\in\mathbb{R}\end{subarray}}\,\frac{1}{n}\sum_{i=1}^{n}\Bigl\{\alpha+\frac{\lambda\beta}{\rho}\log\bigl(1+\rho e^{(\hat{\ell}_{i}(\theta)-\alpha)/(\lambda\beta)}\bigr)\Bigr\},$ (23)
where $\rho>0$ is a parameter controlling the accuracy of the approximation. This approximation can be efficiently optimized with SGD. Note that our method can also be applied to other DRO algorithms such as Sinkhorn DRO (wang2021sinkhorn), which we omit to avoid redundancy.
@@
We consider MNIST dataset (deng2012mnist) with train and validation labels corrupted by feature-dependent noise (see algan2020label) (noise ratio 25%), and original (clean) test labels. Let $\theta$ denote weights of a CNN with two convolutional layers (32 and 64 channels, kernel size 3, ReLU activations, and $\times2$ max pooling), followed by a fully connected classifier with one hidden layer of 128 units, and let $\ell(\theta;z)$ be its cross entropy loss on object $z$. In the experiment, SGD (with batch size 1) is applied to problems (22) (baseline) and (23) (proposed approach). We consider values of the stepsize $\eta\in\{10^{-6},10^{-5},10^{-4},10^{-3}\}$. Parameters $\lambda$ and $\beta$ were set to 1 since smaller values required a smaller stepsize and resulted in a slow convergence, while larger values pushed the model towards fitting the noisy distribution instead of the true one. Approximation accuracy parameter $\rho$ in (23) was set to 0.1. For the inner maximization problem in (21), just 5 iterations of Nesterov’s accelerated gradient method were sufficient to reach plateau in terms of the objective value. Additionally, we used SGD for the usual empirical risk minimization (ERM) to observe the effects of conventional (non-robust) training on noisy data.
Figure 3: Performance of ERM and two DRO approaches on MNIST with noisy labels. Left: ERM accuracy on the noisy validation set vs. clean test set. Middle: validation vs. test accuracy for DRO approaches. Right: training loss $F(\theta)$ from (21).
Figure 3 demonstrates the performance of different approaches with best hyperparameter $\eta$ ($10^{-3}$ for ERM, $10^{-4}$ for the proposed approach, and $10^{-5}$ for the baseline). Shaded regions indicate $\pm$ one standard deviation across 10 runs, except that for the baseline approach (22) we excluded a single run that caused a floating-point exception. The plot on the left illustrates that ERM fits to the corrupted data well (accuracy on the noisy validation set is increasing) which results in decreasing accuracy on the clean test set. In contrast, the plot in the middle shows that an increase in validation accuracy results in the increase in test accuracy for both DRO approaches, which indicates that they are more capable at learning the underlying clean distribution. The plot on the right shows the train loss $F(\theta)$ from (21). As seen from the figure, the proposed approach converges faster than the baseline. This is caused by the fact that the baseline requires a small stepsize to avoid overflows.

### v3

@@ 3.1 Continuous entropy-regularized OT
The classical optimal transport (Monge–Kantorovich) problem consists in finding a coupling of two probablility measures $\mu,\nu\in\mathcal{P}(\mathcal{X})$ which minimizes the integral of a given measurable cost function $c\colon\mathcal{X}\times\mathcal{X}\to\mathbb{R}_{+}$ (e.g., a distance), i.e.,
$W(\mu,\nu)\coloneqq\inf_{\pi\in\Pi(\nu,\mu)}\int c(x,z)\,\mathrm{d}\pi(x,z),$
where $\Pi(\mu,\nu)\subset\mathcal{P}(\mathcal{X}\times\mathcal{X})$ is the set of couplings (transport plans) of $\mu$ and $\nu$ (see Kantorovich, 1942; Villani, 2008; Santambrogio, 2015). For simplicity of demonstration, we assume that the measures are defined on the same space $\mathcal{X}$, but the results extend trivially to the case of two different spaces. Following (Cuturi, 2013), we consider entropy-regularized optimal transport (eOT) problem:
@@
where $\nu\otimes\mu$ is the product measure. It is known that eOT admits the following dual and semi-dual formulations (see, e.g., (Genevay et al., 2016)):
$W_{\varepsilon}(\mu,\nu)$ $=\underbrace{\max_{u,v\in\mathcal{C}(\mathcal{X})}\iint_{{\mathcal{X}\times\mathcal{X}}}f_{\varepsilon}(x,y,u,v)\,\mathrm{d}\mu(x)\,\mathrm{d}\nu(y)}_{\text{dual}}$
$=\underbrace{\max_{v\in\mathcal{C}(\mathcal{X})}\int_{{\mathcal{X}}}h_{\varepsilon}(x,v)\,\mathrm{d}\mu(x)}_{\text{semi-dual}},$ (12)
@@
$f_{\varepsilon}(x,y,u,v)$ $\coloneqq u(x)+v(y)$-
$-\varepsilon\exp\left(\frac{u(x)+v(y)-c(x,y)}{\varepsilon}\right),$ (13)
$h_{\varepsilon}(x,v)$ $\coloneqq\int_{{\mathcal{X}}}v(y)\mathrm{d}\nu(y)-\varepsilon$-
$-\varepsilon\log$ $\left(\int_{{\mathcal{X}}}\exp\left(\frac{v(y)-c(x,y)}{\varepsilon}\right)\mathrm{d}\nu(y)\right),$ (14)
and $\varepsilon>0$ is the regularization coefficient. In the LSOT framework (Seguy et al., 2018), the potentials $u$ and $v$ are parameterized by neural networks and optimized via SGD. While Appendix B.1 contains a more detailed literature review, we briefly position LSOT among other solvers to motivate its selection as a baseline. LSOT offers two key advantages relevant to our goals: it is less computationally intensive than modern solvers requiring adversarial training (Korotin et al., 2023; Gushchin et al., 2023; Asadulaev et al., 2024) or iterative Langevin dynamics (Mokrov et al., 2024), and it supports a general cost function—contrary to other efficient solvers like (Korotin et al., 2024) tailored to the quadratic cost. Therefore, to solve eOT with a general cost function under modest computational constraints, we adopt the LSOT framework as our primary baseline. In Appendix B.2 we compare also to (Genevay et al., 2016) who use an RKHS parametrization for the potentials $u$ and $v$.
@@
The main drawback of this approach is the presence of the exponent in the dual objective (and consequently in the SGD updates). Specifically, exponents are prone to floating-point exceptions (Goldberg, 1991), especially if the regularization parameter $\varepsilon$ is relatively small, which is often the case. For example, if $\varepsilon=0.01$ and $z\geq 7.1$, then $e^{z/\varepsilon}$ exceeds the representable range of a double-precision (float64) floating-point number — an overflow happens. When single precision (float32) is used, an overflow happens even for $z\geq 0.89$.
@@
If we consider instead the semi-dual formulation and use the approximation (6), we get the problem
@@
with
$\tilde{h}_{\varepsilon}(x,y,v,\alpha)\coloneqq v(y)-\alpha(x)-\\ -\frac{\varepsilon}{\rho}\log\bigl(1+\rho e^{(v(y)-c(x,y)-\alpha(x))/\varepsilon}\bigr)-\varepsilon,$ (16)
which also admits neural network parameterization and optimization via SGD. One can show, in the same way as in (Genevay et al., 2016), that this corresponds to the regularized OT problem (11) with Safe KL divergence $D_{\rho}$ rather than the usual KL, i.e.
@@
We consider the MNIST (Deng, 2012) and EMNIST-letters (Cohen et al., 2017) datasets as samples from the distributions $\mu$ (digits) and $\nu$ (letters). Manhattan distance $\ell_{1}$ is chosen as the cost function for computing eOT between $\mu$ and $\nu$. We parameterize the functions $u$, $v$ in LSOT and $v$, $\alpha$ in our proposed approach using a multilayer perceptron with two hidden layers (dimensions 256 and 128) and ReLU activations. The batch size is 256, and the learning rate is selected via grid search over $\{10^{-6},10^{-5},\ldots,10^{-1}\}$. The objective is evaluated on the empirical distributions of the dedicated test sets.
@@
3.2 DRO with KL divergence
Table 1: Objective value (18) (mean $\pm$ std across 10 runs) at epoch 50 for baseline (19) (Levy et al., 2020) and proposed gradient estimator (21) with different $\rho$ values. Results are shown for various penalty coefficients $\lambda$ and batch sizes $|D|$, with optimal learning rates selected from $\{10^{-9},\dots,10^{-4}\}$. Best results per column are shown in bold.
$\lambda=1/5$ $\lambda=1$ $\lambda=5$
One of the approaches to training a model that is robust to data distribution shifts and noisy observations is called Distributionally Robust Optimization (DRO) (Kuhn et al., 2025). In contrast to the standard Empirical Risk Minimization (ERM) approach, which minimizes the average loss on the training sample, DRO minimizes the risk for the worst-case distribution among those close to a reference measure (e.g., empirical distribution). A prominent example is KL divergence DRO (Hu and Hong, 2012), which is formulated as the saddle-point problem
@@ 3.2 DRO with KL divergence
However, when $n$ is large, computing the full gradient $\nabla\mathcal{L}(\theta)=\sum_{i=1}^{n}p_{i}^{*}(\theta)\nabla\ell_{i}(\theta)$ becomes costly. A straightforward approach (Levy et al., 2020) is to sample a batch $D$, compute the respective softmax weights
$p_{i}^{D}(\theta)\coloneqq\frac{e^{\ell_{i}(\theta)/\lambda}}{\sum_{j\in D}e^{\ell_{j}(\theta)/\lambda}},$
and define a gradient estimator by
@@
$\tilde{\nabla}^{D}_{\theta}G(\theta,\alpha)$ $\coloneqq\frac{1}{|D|}\sum_{i\in D}\sigma_{\rho}\left(\textstyle\frac{\ell_{i}(\theta)-\alpha}{\lambda}\right)\nabla\ell_{i}(\theta),$ (21)
$\tilde{\nabla}^{D}_{\alpha}G(\theta,\alpha)$ $\coloneqq 1-\frac{1}{|D|}\sum_{i\in D}\sigma_{\rho}\left(\textstyle\frac{\ell_{i}(\theta)-\alpha}{\lambda}\right).$
@@
Consider the California housing dataset (Pace and Barry, 1997) consisting of 20,640 objects represented by 8 features. Let $\ell_{i}$ be the squared error of a linear model, $\ell_{i}(\theta)=(y_{i}-\theta^{\top}x_{i})^{2}$. We use accelerated SGD with the gradient estimator (19) (Levy et al., 2020) as the baseline approach for solving (18), and compare it to our proposed gradient estimator (21). We consider various values of the penalty coefficient $\lambda\in\{1/5,1,5\}$ and batch sizes $|D|\in\{10,10^{2},10^{3}\}$. For each configuration, we select the optimal learning rate from $\{10^{-9},10^{-8},\ldots,10^{-4}\}$. The approximation accuracy parameter $\rho$ in our method is varied across $\{10^{-1},10^{-3},10^{-5}\}$. Momentum is fixed at 0.9 (without tuning), and the least squares solution is used as the initial point for optimization.
@@
3.3 DRO with unbalanced OT
Table 2: Objective value ($2$) (mean $\pm$ std across 5 runs) at epoch 20 for the baseline (Wang et al., 2024) and the proposed approach (with different $\rho$ values), i.e., SGD on (23) and (24), respectively. Results are shown for various penalty coefficients $\gamma$ and $\lambda$, with optimal learning rates selected from $\{10^{-9},10^{-8},\dots,10\}$. Best results per column are shown in bold.
$\gamma=1/5$ $\gamma=1$ $\gamma=5$
Approach $\lambda=1/5$ $\lambda=1$ $\lambda=5$ $\lambda=1/5$ $\lambda=1$ $\lambda=5$ $\lambda=1/5$ $\lambda=1$ $\lambda=5$
Baseline (23) $1.79\!\pm\!0.01$ $1.78\!\pm\!0.01$ $1.77\!\pm\!0.01$ $0.86\!\pm\!0.01$ $0.81\!\pm\!0.01$ $0.79\!\pm\!0.01$ $0.33\!\pm\!0.06$ $0.22\!\pm\!0.08$ $0.17\!\pm\!0.02$
[…]
(24), $\rho\!=\!10^{-1}$ $1.57\!\pm\!0.16$ $1.29\!\pm\!0.30$ $1.93\!\pm\!0.92$ $\bm{0.45\!\pm\!0.03}$ $\bm{0.37\!\pm\!0.07}$ $\bm{0.31\!\pm\!0.04}$ $\bm{0.21\!\pm\!0.02}$ $0.16\!\pm\!0.09$ $0.12\!\pm\!0.06$
(24), $\rho\!=\!10^{-2}$ $\bm{1.35\!\pm\!0.04}$ $\bm{1.28\!\pm\!0.05}$ $\bm{1.25\!\pm\!0.05}$ $0.50\!\pm\!0.05$ $0.50\!\pm\!0.11$ $0.48\!\pm\!0.02$ $\bm{0.21\!\pm\!0.03}$ $\bm{0.12\!\pm\!0.03}$ $0.26\!\pm\!0.04$
In the KL divergence DRO described in the previous subsection, uncertainty set is limited to distributions with the same support as the empirical measure $\mu=\frac{1}{n}\sum_{i}\delta_{x_{i}}$. Another popular approach, Wasserstein DRO (WDRO) (Mohajerin Esfahani and Kuhn, 2018; Sinha et al., 2020), considers the worst-case risk over shifts within a Wasserstein (OT) ball around a reference measure $\mu$ instead of the KL-ball in (17), thus including continuous probability measures. Unfortunately, this approach is not resilient to outliers that are geometrically far from the clean distribution since OT metric is sensitive to them (Nietert et al., 2023). A natural generalization is to switch to semi-balanced OT (Liero et al., 2018; Chizat et al., 2019; Kondratyev et al., 2016), which replaces a hard constraint on one of the marginals with a mismatch penalty function, e.g.,
$W_{\beta}(\nu,\mu)=\hskip-10.00002pt\inf_{\begin{subarray}{c}\pi\in\mathcal{P}(\mathcal{X}\times\mathcal{X})\\ \pi_{1}=\nu\end{subarray}}\int_{{\mathcal{X}\times\mathcal{X}}}\hskip-10.00002ptc(x,z)\,\mathrm{d}\pi(x,z)+\beta\,D_{KL}(\pi_{2},\mu),$
@@ 3.3 DRO with unbalanced OT
where $\lambda>0$ is the Lagrangian penalty parameter. By standard duality, (Wang et al., 2024) showed that when $\mu=\frac{1}{n}\sum_{i}\delta_{x_{i}}$ is the empirical distribution, this is equivalent to
$\min_{\theta\in\Theta}F(\theta)\coloneqq\gamma\,\log\Bigl(\frac{1}{n}\sum_{i=1}^{n}e^{\hat{\ell}_{i}(\theta)/\gamma}\Bigr)$
$\text{with }\;\hat{\ell}_{i}(\theta)\coloneqq\sup_{z\in\mathcal{X}}\{\ell(\theta;z)-\lambda c(z,x_{i})\},$ (22)
where $\gamma\coloneqq\lambda\beta$. To avoid the costly gradient computation of LogSumExp, the authors drop the logarithm and use SGD to optimize the sum of exponents,
$\min_{\theta\in\Theta}\frac{1}{n}\sum_{i=1}^{n}e^{\hat{\ell}_{i}(\theta)/\gamma}.$ (23)
@@
$\min_{\begin{subarray}{c}\theta\in\Theta\\ \alpha\in\mathbb{R}\end{subarray}}\,\frac{1}{n}\sum_{i=1}^{n}\Bigl\{\alpha+\frac{\gamma}{\rho}\log\bigl(1+\rho e^{(\hat{\ell}_{i}(\theta)-\alpha)/\gamma}\bigr)\Bigr\},$ (24)
where $\rho>0$ is a parameter controlling the accuracy of the approximation. This approximation can be efficiently optimized with SGD. Note that our method can also be applied to other DRO algorithms such as Sinkhorn DRO (Wang et al., 2021), which we omit to avoid redundancy.
@@
We consider MNIST dataset (Deng, 2012) with labels corrupted by feature-dependent noise (see Algan and Ulusoy, 2020) (noise ratio 25%). Let $\theta$ denote weights of a CNN with two convolutional layers (32 and 64 channels, kernel size 3, ReLU activations, and $\times2$ max pooling), followed by a fully connected classifier with one hidden layer of 128 units, and let $\ell(\theta;z)$ be its cross entropy loss on object $z$. In the experiment, SGD with batch size 32 is applied to problems (23) (baseline) and (24) (proposed approach). We consider values of the stepsize $\eta\in\{10^{-9},10^{-8},\ldots,10\}$. For the inner maximization problem in (22), just 5 iterations of Nesterov’s accelerated gradient method were sufficient to reach plateau in terms of the objective value.
Figure 3: Example trajectories of the baseline with $\text{lr}=10^{-4}$ (blue) and the proposed approach with $\rho=0.1$, $\text{lr}=10^{-2}$ (orange). Y-axis corresponds to the objective (22) with $\lambda=\gamma=1$. Green dashed curve illustrates that the baseline diverges during the first epoch even for a relatively small $\text{lr}=10^{-3}$.
Table 2 presents the final objective values (mean $\pm$ standard deviation across 5 runs) after 20 epochs, at which point the methods typically stop making significant progress. The best-performing configuration in each column is highlighted in bold. The table shows that the proposed approach consistently outperforms the baseline. For the approximation parameter $\rho$, a large value ($\rho=1$) often introduces a noticeable approximation gap. An exception occurs when the initial objective is sufficiently smooth, as with $\gamma=5$. The value $\rho=10^{-1}$ offers robust performance across most scenarios, whereas $\rho=10^{-2}$ is better suited for ill-conditioned objectives (e.g., $\gamma=1/5$) that demand high approximation accuracy.
To illustrate the optimization behavior underlying the aggregated results in Table 2, Figure 3 shows example trajectories for a fixed seed and parameter set ($\lambda=\gamma=1$). The proposed approach with $\rho=0.1$ (orange) minimizes the objective using a relatively large learning rate of $10^{-2}$. The baseline (blue), however, requires a learning rate of $10^{-4}$ to avoid numerical overflow, as demonstrated by its rapid divergence (green dashed curve) for $\text{lr}=10^{-3}$. Consequently, the proposed method achieves lower objective values, while the baseline’s small learning rate results in slow progress.